OpenStudy (vbcvbcbcvbvc):

There are nickels and quarters worth $2.20 in total. If there are 28 coins, how many nickels are there?

jimthompson5910 (jim_thompson5910):

Let n = number of nickels q = number of quarters `there are 28 coins` so n+q = 28 Solve for q to get q = 28 - n With me so far?

OpenStudy (vbcvbcbcvbvc):

ok

OpenStudy (vbcvbcbcvbvc):

that's it ?

jimthompson5910 (jim_thompson5910):

`There are nickels and quarters worth $2.20 in total` so we know the other equation to be 0.05*n + 0.25*q = 2.20

OpenStudy (vbcvbcbcvbvc):

ok

jimthompson5910 (jim_thompson5910):

Now plug in q = 28 - n 0.05*n + 0.25*q = 2.20 0.05*n + 0.25*(28 - n) = 2.20 solve for n and tell me what you get

OpenStudy (vbcvbcbcvbvc):

wait ook

OpenStudy (vbcvbcbcvbvc):

24

jimthompson5910 (jim_thompson5910):

good, now use this to find q q = 28 - n plug in n = 24

OpenStudy (vbcvbcbcvbvc):

4?

jimthompson5910 (jim_thompson5910):

n = 24 q = 4 so there are 24 nickels and 4 quarters

OpenStudy (vbcvbcbcvbvc):

my answer is 24 or 4 ?

jimthompson5910 (jim_thompson5910):

Go back to what it's asking `If there are 28 coins, how many nickels are there?`

OpenStudy (vbcvbcbcvbvc):

oh 24

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (vbcvbcbcvbvc):

Thank you, You are a life saver

OpenStudy (vbcvbcbcvbvc):

i am posting another question